# Fair and Truthful Allocations Under Leveled Valuations

### v1

@@ 5 Truthfulness
[Serial dictatorship with quotas, Pápai (2000)] The serial dictatorship with quotas (SDQ) mechanism is specified by a permutation of the agents $\sigma:N\rightarrow N$ and a quota system $q:N\rightarrow N$, $q=(q_{1},\dots,q_{n})$ such that $\sum_{i=1}^{n}q_{i}=m$.
5.1 Maximin Share Guarantee
In the case of additive agents, it is known that an EFX allocation is also a $2/3$-MMS allocation for $n\in\{2,3\}$, due to Amanatidis et al. (2018). The approximation degrades as the number of agents grow to $4/7=0.5714$. We show that any allocation obtained via SDQs with the quota system of Algorithm 1 achieves a $2/3$-MMS approximation (along with EFX), even for $n>3$.
@@ 5.1 Maximin Share Guarantee
There exists a $\frac{2}{3}$-MMS allocation under additive leveled valuations.
@@
Consider an arbitrary allocation $B=(B_{1},\dots,B_{n})$ obtained via Algorithm 1. Let $i$ be an agent that receives $B_{i},\lvert B_{i}\rvert=k=\lfloor m/n\rfloor$ and let allocation $B^{\prime}=(B_{1}^{\prime},\dots,B_{n}^{\prime})$ denote her MMS allocation. Assume wlog that $k\geq 2$; the case where $m<2n$ is handled by Proposition 3.1 via an SDQ. Note that since the valuations are additive, there must exist an item $g\in B_{i}$ such that $v_{i}(g)\leq v_{i}(B_{i})/2$ since $k\geq 2$. Otherwise, the items would add up to something larger than $v_{i}(B_{i})$. If all agents’ bundles are of size $k$, then $g$ belongs to a bundle, say $B_{j}^{\prime}$, in her MMS allocation $B^{\prime}$. We know that $\lvert B_{j}^{\prime}\setminus g\rvert=k-1$, and thus, due to leveled valuations, $v_{i}(B_{j}^{\prime}\setminus g)\leq v_{i}(B_{i})$. Therefore, we have $\mu_{i}\leq v_{i}(B_{i})+v_{i}(g)\leq 3/2v_{i}(B_{i})$. Differently, some agents may receive one less item than others as shown in Algorithm 1. But then, $\mu_{i}<v_{i}(B_{j}^{\prime})$ for some bundle $B_{j}^{\prime}$ with $\lvert B_{j}^{\prime}\rvert=k+1$ that contains $g$, and thus, $v_{i}(B_{j}^{\prime}\setminus g)\leq v_{i}(B_{i})$ (since $i$ precedes $j$ in the order). We conclude that $\mu_{i}\leq v_{i}(B_{j}^{\prime})\leq v_{i}(B_{i})+v_{i}(g)\leq 3/2v_{i}(B_{i})$. ∎
@@
Any balanced allocation achieves a $\frac{k-1}{k}$-MMS approximation under additive leveled valuations, where $k=\lfloor m/n\rfloor$.
@@
Let $k=\lfloor m/n\rfloor$ and let $B_{i}^{\prime}$ be the bundle that gives the MMS value to agent $i$, i.e., $v_{i}(B_{i}^{\prime})=\mu_{i}$. In any balanced allocation (arbitrary SDQ following the quota system of Algorithm 1), an agent $i$ either receives $k+1$ items (in which case she trivially guarantees her MMS) or exactly $k$ items. We examine the latter case. We know that the MMS bundle consists of $k$ items, and thus, any bundle $B$ with $\lvert B\rvert=k$ has higher value than any subset $S_{j}$ of $B_{i}^{\prime}$ with cardinality $\lvert S_{j}\rvert=k-1$; there are $\binom{k}{k-1}=k$ such combinations (leaving out one item at a time). We derive a set of equations $v_{i}(B)\geq v_{i}(S_{j})$ for all $j\in[k]$. Summing over all inequalities, we get $k-1$ copies of $B_{i}^{\prime}$ in the right-hand side. Hence, we have: $k\cdot v_{i}(B)\geq(k-1)\cdot\mu_{i}\implies v_{i}(B)\geq\frac{k-1}{k}\cdot\mu_{i}$
It is easy to see that the aforementioned propositions imply that fixed-point mechanisms and SDQs trivially achieve strong approximation properties. Lastly, one can derive analogous propositions for truthful MMS approximations via SDQs for subadditive valuations.
Proposition 5.3.
There exists a $\frac{1}{2}$-MMS approximation under subadditive leveled valuations via an SDQ.
5.2 Envy-freeness up to Any Good
We focus on (strict) ordinal preferences. We give an algorithmic characterization of the mechanisms satisfying a set of desired properties. The same characterization holds for the case of lexicographic valuations Hosseini et al. (2021).
Proposition 5.4.
(Pápai (2000)) For leveled preferences, a mechanism is Pareto optimal, strategyproof, non-bossy, and neutral if and only if it is an SDQ.
Theorem 5.
A mechanism satisfies EFX, PO, truthfulness, non-bossiness, and neutrality if and only if it is a Serial Quota Dictatorship.
Proof.
Algorithm 1 is an SDQ with quotas $q_{i}=\lfloor m/n\rfloor$ for all $i\in[n-r]$ and $q_{i}=\lfloor m/n\rfloor+1$ for all $i>n-r$, and therefore EFX according to Theorem 4. Moreover, it follows from Proposition 5.4 that it is PO, truthful, non-bossy, and neutral. We also prove the converse; assume that there exists a mechanism $f$ that satisfies these properties. By Proposition 5.4, $f$ must be an SDQ for some $\sigma$ and a specified quota system such that $\sum_{i=1}^{n}q_{i}=m$. Assume wlog that $m\geq n$. Suppose, for the sake of contradiction, that the quota system of $f$ violates the quota system described by Algorithm 1; then there exists a pair of agents $i,j\in N$ such that $\lvert A_{i}\rvert=\lvert A_{j}\rvert-2$, violating EFX due to the definition of leveled valuations, which is a contradiction. Therefore, $f$ must be an SDQ with the quota system specified in Algorithm 1. ∎
@@ 5.2 Envy-freeness up to Any Good
Any deterministic strategyproof and non-bossy mechanism is also group-strategyproof Pápai (2000).

### v2

@@ 5.1 Maximin Share Guarantee
[Serial dictatorship with quotas, (Pápai, 2000)] The serial dictatorship with quotas (SDQ) mechanism is specified by a permutation of the agents $\sigma:N\rightarrow N$ and a quota system $q=(q_{1},\dots,q_{n})$ such that $\sum_{i=1}^{n}q_{i}=m$.
Corollary 4.
There exists a deterministic truthful mechanism that guarantees EFX and $\frac{k-1}{k}$-MMS for additive leveled valuations, where $k=\lfloor m/n\rfloor$ and $k\geq 2$.
Lastly, one can derive analogous propositions for truthful MMS approximations via SDQs for subadditive valuations.
@@
An $\frac{1}{2}$-MMS approximation under subadditive leveled valuations can be achieved via a truthful mechanism.
@@
Consider an SDQ mechanism with the quota structure of Algorithm 1. We have argued that the bundle that corresponds to the MMS value has cardinality $k$. Let’s consider the case of $k\geq 2$; the case of $k<2$ is treated by Proposition 3.1 which can be used with an SDQ. Let’s assume that there is an agent $i$ receiving a bundle $B_{i}$ that does not achieve her maximin share, i.e., $v_{i}(B_{i})\leq\mu_{i}$, otherwise the claim follows. Let $g\not\in B_{i}$ be an arbitrary item. Then, we have:
$\mu_{i}\leq v_{i}(B_{i}\cup g)\leq v_{i}(B_{i})+v_{i}(g)\leq 2\cdot v_{i}(B_{i})$
and the claim follows. The inequalities follow from subadditivity and the definition of leveled valuations. ∎
Corollary 5.
There exists a deterministic truthful mechanism that simultaneously guarantees EFX and $\frac{1}{2}$-MMS for subadditive leveled valuations.
Under leveled valuations, we observe that strategyproofness can be achieved in conjunction with strong fairness guarantees. Amanatidis et al. (2017a) introduce the notion of controlling items; we say that a player controls an item if reporting certain values guarantees her this item. Control is the major hurdle that makes fairness and truthfulness incompatible under additive valuations. We note that this is not applicable to our case, since control of singletons [^7] Control of singletons implies that whenever an agent strongly desires an item, i.e. prefers it more than all the other items combined, she gets it. (or pairs) is not possible under leveled valuations.
5.2 Envy-freeness up to Any Good
We focus on (strict) ordinal preferences. We give an algorithmic characterization of the mechanisms satisfying a set of desired properties. The same characterization holds for the case of lexicographic valuations (Hosseini et al., 2021).
Axiomatic Properties. We are interested in core axiomatic properties, such as efficiency, non-bossiness, and neutrality. A common notion of economic efficiency is Pareto optimality. Roughly speaking, under a Pareto efficient allocation, no individual can become strictly better off without hurting someone else’s value. There exist instances where EFX cannot be satisfied in conjunction with Pareto optimality, even for 2 players (Plaut and Roughgarden, 2020). Non-bossiness is a property of the allocation rule that ensures that no agent can change (by a deviating bid) the allocation without changing their own bundle. Neutrality states that the allocation rule does not depend on the labeling of the items.
@@ 5.2 Envy-freeness up to Any Good
(Pápai, 2000) For leveled preferences, a mechanism is Pareto optimal, strategyproof, non-bossy, and neutral if and only if it is an SDQ.
Theorem 5.
A mechanism satisfies EFX, PO, truthfulness, non-bossiness, and neutrality if and only if it is a Serial Quota Dictatorship with the quotas of Algorithm 1.
@@
Algorithm 1 is an SDQ with quotas $q_{i}=\lfloor m/n\rfloor$ for all $i\in[n-r]$ and $q_{i}=\lfloor m/n\rfloor+1$ for all $i>n-r$, and therefore EFX according to Theorem 4. Moreover, it follows from Proposition 5.3 that it is PO, truthful, non-bossy, and neutral. We also prove the converse; assume that there exists a mechanism $f$ that satisfies these properties. By Proposition 5.3, $f$ must be an SDQ for some $\sigma$ and a specified quota system such that $\sum_{i=1}^{n}q_{i}=m$. Assume w.l.o.g. that $m\geq n$. Suppose, for the sake of contradiction, that the quota system of $f$ violates the quota system described by Algorithm 1; then, we show that EFX is violated. First, note that if there exists a pair of agents $i,j\in N$ such that $\lvert A_{i}\rvert=\lvert A_{j}\rvert-2$, then EFX is violated, as agent $i$ will envy agent $j$ due to the definition of leveled valuations, which is a contradiction. So it must be that the sizes of the allocated bundles must differ by at most one. Then we argue that if there is an agent that receives a smaller size bundle, they must precede those that receive the larger size bundles. Indeed, suppose that agent $j$ picks her bundle before agent $i$, i.e., $\sigma(j)<\sigma(i)$, but $\lvert A_{j}\rvert=\lvert A_{i}\rvert+1$. Then, it may be the case that $j$ has picked $i$’s favorite available bundle of size $\lvert A_{i}\rvert$ along with some additional item $g$, thus violating EFX after the hypothetical removal of $g$. Therefore, we conclude that $f$ must be an SDQ with the quota system specified in Algorithm 1. ∎
@@
Any deterministic strategyproof and non-bossy mechanism is also group-strategyproof (Pápai, 2000).